OpenStudy (anonymous):

Whole milk contains 3.7% fat. You want to add 2%-fat milk to 5 fI oz of whole milk. How many fl oz 0f 2% milk must you add?

OpenStudy (anonymous):

to make what?

OpenStudy (anonymous):

Oh sorry, to make 3%-fat milk :)

OpenStudy (anonymous):

5*0,037 + x*0,02 = (5+x) * 0,03

OpenStudy (anonymous):

5*0,007 = x*0,01 x= 5*0,007/0,01